OpenStudy (anonymous):

39x^7y^5/3x^3y simplify it

OpenStudy (amistre64):

\[\frac{39}{3}\frac{x^7}{x^5}\frac{y^5}{y}\] should make it easier to see what has to happen

OpenStudy (amistre64):

but then i spose it helps if you put the right exponents on the appropriate spots eh

OpenStudy (amistre64):

its spose to be an x^3 under there

OpenStudy (anonymous):

oh it gets confusing when u get to the y part what r u supposed to do?

OpenStudy (amistre64):

y^5/y ? yyyyy ---- = yyyy y

OpenStudy (anonymous):

y^4?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

the asnswer is 13x^4y^4

OpenStudy (amistre64):

looks good to me

OpenStudy (anonymous):

ok thanks